Find all 4 of the second order partial derivatives for the function f(x,y)=e^-xy +x^2y cos(x) -y^2


1
Expert's answer
2021-09-30T09:35:58-0400

"f_x=-ye^{-xy}+2xycosx-x^2ysinx"

"f_y=-xe^{-xy}+x^2cosx-2y"

"f_{xx}=y^2e^{-xy}+2ycosx-2xysinx-2xysinx-x^2ycosx="

"=y^2e^{-xy}+ycosx(2-x^2)-4xysinx"

"f_{xy}=f_{yx}=-e^{-xy}+xye^{-xy}+2xcosx-x^2sinx="

"=e^{-xy}(xy-1)+2xcosx-x^2sinx"

"f_{yy}=x^2e^{-xy}-2"